### v2

# Revealing the nature of ultra-long period objects with space-based gravitational-wave interferometers

Abstract
A few members of the recently-discovered class of ultra-long period objects have been identified as binaries with white-dwarf primaries. In most cases however, electromagnetic data are inconclusive and isolated magnetars or compact binaries remain viable. If the pulsation period matches that of the orbit though – as is the case for ILT J1101+5521 and GLEAM-X J0704–37 – some of these elusive radio transients could be gravitational-wave bright in the mHz band. Space-based interferometers could thus be used to provide independent constraints on their nature. We quantify the signal-to-noise ratio for the known systems, under a variety of scenarios, and show that a few could be detectable for sufficiently large chirp masses. Astrophysical implications for (non-)detections are discussed.
gravitational waves, binaries: close, stars: white dwarfs, neutron stars
[…]
=3
1 Introduction
Ultra-long period objects (ULPs) are an elusive class of coherent, radio-pulsating transients with periods of $P\sim$ hours that have thus far evaded classification. Their brightness variations imply compact sources and, given spectral similarities to pulsar emissions, the natural interpretation is that they are slow neutron stars (see Coti Zelati & Borghese, 2024, for a discussion). However, this is theoretically challenging to accept as their radio luminosities ($L_{\nu}$) categorically exceed the implied spindown luminosity, necessitating an alternative power reservoir. While magnetic fields could provide such a battery for magnetar progenitors (Beniamini, Wadiasingh & Metzger, 2020; Rea et al., 2022; Cooper & Wadiasingh, 2024), optical spectroscopy has confirmed that some ULPs are binaries consisting of white dwarfs with M-dwarf companions (WDMDs; Hurley-Walker et al., 2024; de Ruiter et al., 2025). As observational efforts are revealing a sizeable population of ULPs, residing within new portions of the $P$-$\dot{P}$, $L_{\nu}$-$L_{\rm X}$, and other parameter spaces, tools to constrain their nature are of astrophysical relevance.
Because pulsations from ILT J1101+5521 and GLEAM-X J0704–37 – confirmed WDMDs – are phase-aligned with their orbital motion (de Ruiter et al., 2025; Rodriguez, 2025), this may be the case for other ULPs. On the other hand, the lowest-period systems cannot be WDMDs with $P\approx P_{\text{orb}}$ because the companion would not fit inside its Roche lobe. Even so, because the maser-like radio emission mechanisms often invoked to explain pulsations in binaries do not rely on companion makeup (Melrose, 2017; Qu & Zhang, 2025) a pair of compact objects, interacting magnetically in a tight orbit, are also viable progenitors (Lyman et al., 2025).
We point out that because several ULPs reside at sub-kpc distances, they may be strong sources of gravitational waves (GWs) independently of the radio mechanism if the orbital period corresponds to the pulsation period. As eccentricity should be erased before a binary compactifies to $P_{\text{orb}}\sim$ hours (Peters, 1964), emissions at frequencies of $f_{\rm GW}=2/P_{\text{orb}}\approx 0.6\,\text{mHz}\times\left({1\,\text{hr}}/{P_{\text{orb}}}\right)$ ought to dominate the spectrum in this scenario. Within the next decade, a network of space-based, GW interferometers will become operational, namely the Laser Interferometer Space Antenna (LISA; Amaro-Seoane et al., 2017) plus the two Chinese-led projects Taiji (Luo et al., 2020) and TianQin (Huang et al., 2020). These detectors will be most sensitive in the $\gtrsim$ mHz band and can search for signals from ULPs. Consequently, future GW observations can be used to effectively confirm or rule out binary scenarios with two heavy objects in several systems where electromagnetic data are inconclusive.
This short paper is organised as follows. Multiwavelength observations and binary scenarios more generally are reviewed in Section 2 to pave the way for estimates of the characteristic strains from ULPs in Section 3. Signal-to-noise ratios (SNRs) are quantified in Section 3.1, where it shown that several ULPs are promising candidates for GW detection under favourable conditions. Some discussion on astrophysical implications are also provided, with conclusions given in Section 4.
2 Constraints on binary nature
Given their observed brightness temperatures, compact sources are required for ULPs to emit coherently (e.g. Qu & Zhang, 2025). With this in mind, the basic question we pose is that if ULPs are (i) composed of a binary involving at least one compact object and (ii) pulsating at the orbital period, would they be visible to space-based GW interferometers? From a detection standpoint, the main ingredients needed to assess this are the distance, $d$, and the chirp mass,
$\mathcal{M}=\frac{\left(M_{\star}M_{\rm c}\right)^{3/5}}{\left(M_{\star}+M_{\rm c}\right)^{1/5}},$ (1)
for primary and companion masses $M_{\star}$ and $M_{\rm c}$, respectively. While ULP distances are reasonably constrained thanks to their dispersion measures, system masses (or indeed binary nature) are much more uncertain in most cases. In order to estimate the viable chirp-mass range and the viability of the hypothesis generally, we begin by reviewing the main constraints in this context.
2.1 White-dwarf plus M-dwarf
Two ULPs—ILT J1101 and GX J0740—have been confirmed as WDMDs with $P=P_{\rm orb}$ from broadband photometry, optical spectroscopy, and other channels (Hurley-Walker et al., 2024; de Ruiter et al., 2025; Rodriguez, 2025). Their data respectively indicate $M_{\rm c}\approx 0.19M_{\odot}$ and $M_{\rm c}\approx 0.32M_{\odot}$ (though cf. Rodriguez, 2025, for a Monte Carlo analysis). While primary masses cannot be dynamically-inferred with certainty due to the unknown inclination, the discovery papers suggest $M_{\star}\sim 0.6M_{\odot}$ and $M_{\star}\sim 0.8M_{\odot}$, respectively.
For ILT J1101 however, de Ruiter et al. (2025) note that arguments based on Roche-lobe overflow indicate the tighter bound [^1] This restriction applies for an M-dwarf radius of $R_{\rm MD}=0.217R_{\odot}$. If the radius was lower by even a few percent ($R_{\rm WD}\lesssim 0.21R_{\odot}$), the primary mass becomes practically unconstrained (see their extended data figure 8). $M_{\star}\lesssim 0.3M_{\odot}$. In particular, observations of M dwarfs indicate a mass-radius relationship of $M_{\rm MD}/M_{\odot}\approx R_{\rm MD}/R_{\odot}$ (Parsons et al., 2018), with the lightest (confirmed) object being EBLM J0555–57AB with $M_{\rm MD}=0.084(4)M_{\odot}$ (von Boetticher et al., 2019). Using the Eggleton (1983) formula and assuming a Keplerian orbit, the radius of the Roche sphere reads
$R_{\rm R}$ $\approx a\times\frac{0.49q^{2/3}}{0.6q^{2/3}+\log(1+q^{1/3})}$ (2)
$\gtrsim 0.1R_{\odot}\left(\frac{P_{\rm orb}}{60\text{ min}}\right)^{2/3},$
for orbital separation $a$ and mass-ratio $q$. Since $R_{\rm MD}\leq R_{\rm R}$ is required for stability, an effective upper limit to $M_{\star}$ via $q$ is set by expression (2) if $P=P_{\rm orb}$. Such an argument further implies that it is unlikely that ULPs with $P\lesssim 30$ min are WDMDs pulsing at the orbital period (see Table 1 for our considered sample). Alternative scenarios with either more compact progenitors, $P\neq P_{\rm orb}$, or isolated stars are required in these cases.
Table 1: Pulsation periods (in descending order) and dispersion-measure distances for the ULPs considered in this work. Some source names have been shortened.
Source name Period (s) Distance (kpc)
[…]
Notes. ${}^{\text{a}}$ Dong et al. (2024) ${}^{\text{b}}$ Hurley-Walker et al. (2022) ${}^{\text{c}}$ Hurley-Walker et al. (2023) ${}^{\text{d}}$ Wang et al. (2025) ${}^{\text{e}}$ Caleb et al. (2024) ${}^{\text{f}}$ Hyman et al. (2009) ${}^{\text{g}}$ de Ruiter et al. (2025) ${}^{\text{h}}$ Hurley-Walker et al. (2024) ${}^{\text{i}}$ Lee et al. (2025).
With respect to the possibility that $P\neq P_{\rm orb}$, there are two known white-dwarf pulsars (J1912 and Ar Sco) with periods that do not correspond to the orbit but rather the spin of the primary (Marsh et al., 2016; Pelisoli et al., 2023, 2024). However, the radio profiles of these objects are notably different from the two confirmed WDMD ULPs. Their observed (highest peak) radio luminosities, $\sim 2\times 10^{26}$ erg/s at $\approx 1$ GHz, are much lower than the values of ILT J1101 ($\sim 1.5\times 10^{28}$ erg/s at $\approx$1 GHz) and GX J0740 ($\sim 10^{28}$ erg/s at $\approx$100 MHz); see appendix B.3 of Rodriguez (2025). Additionally, the degree of linear polarization in J1912 and Ar Sco are at the level of a few percent while the ULPs ILT J1101 and GX J0740 display values of at least $\approx 20\%$ (see table 1 in Qu & Zhang, 2025). Ar Sco and J1912 also shine persistently in X-rays, while ULPs do not. The operative pulsation mechanisms and binary characteristics are likely different therefore between these astronomical classes; we exclude J1912 and Ar Sco from our sample for these reasons.
For WDMD scenarios, we take a modest value of
$\mathcal{M}_{\rm WDMD}\approx 0.3M_{\odot},$ (3)
consistent with the mid-range obtained from the orbital solutions for ILT J1101 ($\mathcal{M}\approx 0.29M_{\odot}$) and GX J0704 ($\mathcal{M}\approx 0.43M_{\odot}$) quoted earlier.
2.2 White dwarfs with compact companions
If the shortest period ULPs are not WDMDs, what type of systems could be they be? While isolated neutron stars are viable – and perhaps even favoured in most cases (Beniamini, Wadiasingh & Metzger, 2020; Cooper & Wadiasingh, 2024) – double degenerates are harder to rule out owing to their intrinsic dimness (see Lyman et al., 2025, for a discussion).
[…]
In scenarios of a double dwarf (or dwarfs with a compact companion which we do not distinguish notationally), we consider a (optimistic) chirp mass of
$\mathcal{M}_{\rm WDWD}\approx 0.5M_{\odot}.$ (4)
The value (4) is slightly higher than the range inferred for the tight AM CVns HM Cancri ($\mathcal{M}\approx 0.45M_{\odot}$), V803 Cen ($\mathcal{M}\lesssim 0.26M_{\odot}$), and V407 Vul ($\mathcal{M}\lesssim 0.35M_{\odot}$; see table 2 in Solheim, 2010), though this may be justified by noting that dwarf masses are correlated with magnetic field strengths (Kawka, 2020) and strong fields may be required to instigate radio pulsing. For a more conservative estimate, the reader can instead examine our results for the chirp mass from expression (3) even with a WDWD picture (see Sec. 3).
2.2.1 CHIME J0630+25
[…]
$\dot{P}_{\rm GW}\approx-5\times 10^{-11}\left(\frac{P_{\rm orb}}{421\text{ s}}\right)^{-\tfrac{5}{3}}\left(\frac{\mathcal{M}}{0.5M_{\odot}}\right)^{\tfrac{5}{3}}\text{ss}^{-1}.$ (5)
Secondly, the lack of an optical counterpart is concerning owing to the closeness of the source ($d\sim 100$ pc), as many WDWDs are brighter (e.g. ZTF J153932.16+502738.8 with an orbital period of $\sim 7$ min is optically visible despite its $\sim 2.3$ kpc distance; Burdge et al., 2019). Dust obscuration or binary inclination could alleviate this tension to some degree, though the above pose a challenge for a WDWD interpretation. On the other hand, the fact that the best-fit $\dot{P}$ is negative hints at a binary, as emphasised by Dong et al. (2024).
All considered, we fix $\mathcal{M}_{\textrm{J}0630}=0.1M_{\odot}$ to avoid tension with expression (5), remarking that orbital decay could be accelerated by tides or electromagnetic interactions (decreasing the permitted $\mathcal{M}$) or decelerated if dissipative torques transfer angular momentum from the primary to the orbit (increasing it; see Marsh, Nelemans & Steeghs, 2004). We emphasise however that this source, together with DA J1832 (see Wang et al., 2025; Suvorov, Dehman & Pons, 2025), is probably the least likely candidate for a binary ULP. Nevertheless, the point we make in this work is that LISA can constrain binary scenarios in a way that electromagnetic data cannot: even if $\mathcal{M}$ was lower by an order of magnitude, the closeness of this source together with its short period would permit detection if $P=P_{\rm orb}$ (Sec. 3.1).
2.2.2 GX J1627 and GPM J1839
[…]
GPM J1839 stands out amongst ULPs as this source has displayed a high duty cycle ($\sim 25\%$) going back to 1988 (Hurley-Walker et al., 2023). In a neutron-star scenario, this points towards aggressive crustal activity despite being old and ‘cold’ (Cooper & Wadiasingh, 2024), which could be used to implicitly constrain magnetar evolution (e.g. Dehman et al., 2020). This is supported by observations of linear-to-circular polarisation conversion taking place in a manner remarkably reminiscent of the transient magnetar XTE J1810–197 (Men et al., 2025). If instead confirmed as a binary, independently of whether $P=P_{\rm orb}$, the fact that the radio luminosity exceeds that of AR Sco by several orders of magnitude indicates that pulsations from binaries involving dwarfs could be highly variable, important to account for in future radio surveys. Dedicated X-ray monitoring would also be useful to constrain source nature as the current upper limit, $L^{\rm max}_{\rm X}\sim 10^{33}$ erg/s, is relatively weak.
2.3 Neutron stars plus one
Figure 1: Characteristic strains, equation (8), for ULPs listed in Tab. 1 (see legends) assuming $P=P_{\text{orb}}$ and $T=2$ yr. Objects are colour-coded according to different scenarios corresponding to chirp masses $\mathcal{M}$ set by expressions (3) (green), (4) (red), or (6) (blue). Two exceptions are CHIME J0630 (Sec. 2.2.1) and GPM J1839 (Sec. 2.2.2), which could be double-compact systems but where orbital braking maxima restrict the component masses if there is no active mass transfer. Filled symbols correspond to cases which exceed or overlap with the LISA strain curve, taken from data provided by Robson, Cornish & Liu (2019).
Except where the above exclusions apply, the chirp mass could be considerably higher if invoking a compact neutron-star/white-dwarf binary like PSR J1141–6545 ($P_{\rm orb}\approx 4.7$ hr and $\mathcal{M}\sim 1\,M_{\odot}$; Bailes et al., 2003) or a neutron/helium star binary like PSR J1928+1815 ($P_{\rm orb}\approx 3.6$ hr and $\mathcal{M}\sim 1.2M_{\odot}$; Yang et al., 2025). Binaries involving mature neutron stars are particularly difficult to constrain with multiwavelength data, much like isolated objects. While it is more plausible that the pulsation periods correspond to the rotation rate of the primary in such cases – and indeed it has been argued that mass-loaded winds, magnetospheric twists, or interactions with a fallback disk could allow magnetars to reach ultra-long periods (Beniamini, Wadiasingh & Metzger, 2020; Beniamini et al., 2023; Suvorov, Dehman & Pons, 2025) – the assumption $P=P_{\rm orb}$ is difficult to exclude a priori. The final (most optimistic) scenario we consider corresponds to a J1141-like chirp mass of
$\mathcal{M}_{\rm NS}\approx 1M_{\odot},$ (6)
where the subscript indicates a degree of agnosticism with respect to a compact companion. For example, Turolla, Possenti & Treves (2005) suggested that GCRT J1745 could contain a double neutron-star system with $\mathcal{M}\approx 1.13M_{\odot}$ and $P=P_{\rm orb}$ where radio emission originates from shocks within the interaction zone of the primary’s wind and the companion’s magnetosphere.
We close this section by stressing that we have not attempted to present a complete survey of ULP scenarios, as these can be found in the respective discovery papers and literature (see, e.g., Rea et al., 2022; Coti Zelati & Borghese, 2024). Indeed, our goal here concerns GW visibility, for which the exact binary makeup is unimportant (though is, of course, important with respect to implications of detections). Even if certain types of binarity may be disfavoured in individual cases, we consider a range of chirp masses to estimate SNRs. Larger or smaller $\mathcal{M}$ values can be easily substituted using the estimates provided in the following sections.
3 Orbital gravitational waves
The intrinsic GW amplitude of a circular binary reads
$h_{0}\approx 2\times 10^{-22}\left(\frac{1\text{ hr}}{P_{\text{orb}}}\right)^{2/3}\left(\frac{\mathcal{M}}{0.5M_{\odot}}\right)^{5/3}\left(\frac{1\text{ kpc}}{d}\right).$ (7)
Assuming that the pulse period is that of the orbit, we use the data collated in Table 1 to estimate $h_{0}$ from equation (7) for each ULP using the chirp mass spread considered in Sec. 2. Notably, the period derivative is moderate in most systems ($|\dot{P}|_{\rm max}\lesssim 10^{-9}\text{ ss}^{-1}$; Coti Zelati & Borghese, 2024), meaning that GW signals should be essentially monochromatic over multi-year observational windows, $T$. This allows for an accumulation of signal power over many orbital cycles, $N=Tf_{\rm GW}=2T/P_{\rm orb}$. Averaging over orbital orientations and polarisations, the characteristic strain $h_{c}$ felt by a detector can thus be estimated through $h_{c}\approx\sqrt{2N}h_{0}$ (Finn & Thorne, 2000), with value
$h_{c}\approx$   $\,\,2.1\times 10^{-20}\left(\frac{1\text{ hr}}{P_{\text{orb}}}\right)^{7/6}\left(\frac{\mathcal{M}}{0.5M_{\odot}}\right)^{5/3}$ (8)
×$\qquad\times\left(\frac{T}{0.5\text{ yr}}\right)^{1/2}\left(\frac{1\text{ kpc}}{d}\right).$
Using the LISA design specifications from Amaro-Seoane et al. (2017), we adopt the noise power spectral density ($S_{n}$) fits from Robson, Cornish & Liu (2019), which include intrinsic noise (from single-link optical metrology and test mass acceleration) as well as Galactic confusion noise from unresolved sources, to estimate ULP detectability. Figure 1 shows characteristic strains $h_{c}$ from equation (8), as functions of GW frequency, for objects listed in Tab. 1 assuming an observation time of two years. We find that two of the known ULPs (solid symbols) could be detectable by LISA (and also by Taiji and TianQin, though not shown), with a couple others being marginal, under favourable conditions (see Sec. 3.1). As it happens, the most promising candidates are also among the most intriguing.
CHIME J0630. Given its proximity to Earth, insights on local supernovae rates could be gleaned if this source is a neutron star, a position which would be firm if no detection is made shortly after LISA launch. The period would thus correspond to rotation, requiring that some magnetars spindown faster than stipulated by pure dipole braking (e.g. via mass-loaded winds; Beniamini, Wadiasingh & Metzger, 2020). An excess of nearby neutron stars (and magnetars in particular) could also call for revisions to population synthesis models (though see Popov et al., 2003). While a nearby white dwarf is not particularly unusual, Tauris (2018) estimates that for SNRs $\gtrsim 100$, $\mathcal{M}$ could be measured to within $\lesssim 1\%$ from which constraints on binary makeup, and possibly even pre-Newtonian effects predicted by some modified theories of gravity (Littenberg & Yunes, 2019), follow.
GX J1627. Discovered in 2018 archival data from the Murchison Widefield Array, this object has since been silent (Hurley-Walker et al., 2022). If interpreted as a neutron star, because the spindown luminosity, $L_{\rm sd}\lesssim 10^{28}$ erg/s, is much lower than the radio luminosity, $L^{1.4\text{GHz}}_{\nu}\sim 4\times 10^{31}$ erg/s, it could not be rotation powered. If magnetically powered, the deep X-ray limits set by Chandra ($L^{\rm max}_{\rm X}\sim 10^{30}$ erg/s; Rea et al., 2022) mean that a non-detection by LISA could indicate that (some) ULPs populate a magnetar branch that is distinct from soft gamma repeaters and anomalous X-ray pulsars (Suvorov, Dehman & Pons, 2025). For a dwarf, the lack of radio reactivation points towards either a choked magnetosphere (e.g. via accretion) or that emissions are beamed inconsistently, either of which holds promise to inform on tight-binary interaction.
DA J1832. Because this ULP has shown contemporaneous X-ray bursting with radio pulsing (Wang et al., 2025), a magnetar progenitor is favoured over a binary (Suvorov, Dehman & Pons, 2025). If confirmed as a binary however, sporadic accretion could be responsible for igniting a burst of comparable luminosity to that seen from DA J1832 in February 2024 ($L_{\rm X}\sim 10^{33}$ erg/s), though it would be difficult to explain the lack of any quiescent emissions (see Mukai, 2017). If the object enters into an outburst phase in future, searches for absorption lines could be useful to constrain its nature. The fact that the source was radio-loud for a $\sim$ year is also uncharacteristic for transitional pulsars (for instance), and so a detection could insist on revisions to theoretical models from a variety of directions.
A J1935. This source displays a unique property amongst ULPs, being that of state switching: there are pulse states with high-degrees of linear polarisation, others which are predominantly circular, and a null state (Caleb et al., 2024). A high degree of linear polarisation and frequent nulling is rather reminiscent of radio-loud magnetars. For an isolated object, a highly-dynamic magnetosphere would be required to explain this temperamental behaviour; the evolution could be driven, for example, by helicity injections from the plastic motions of crustal platelets (e.g. Beloborodov, 2009).
3.1 Signal-to-noise ratios
In general, a confident detection requires an SNR of at least $\sim$ 10. While detection odds improve for sources with known sky locations, we provide conservative estimates by avoiding complications related to barycentric and other corrections. Borrowing formula (26) from Robson, Cornish & Liu (2019), the orientation- and polarisation-averaged SNR is estimable through
$\text{SNR}\approx\frac{8\pi^{2/3}G^{5/3}\mathcal{M}^{5/3}f_{\rm GW}^{2/3}}{\sqrt{5}c^{4}d}\sqrt{\frac{T}{S_{n}(f_{\rm GW})}}\ ,$ (9)
for speed of light $c$ and Newton’s constant $G$. We calculate expression (9) for sources listed in Tab. 1. The results are shown in Table 2 for two scenarios: (i) using the maximum DM-inferred distances and lowest values of $\mathcal{M}$ considered in Sec. 2, and (ii) an optimistic case where the masses and distances take their upper and lower confidence bounds, respectively. The latter scenario is indicated by bracketed values in the second ($T=0.5$ yr) and third ($T=4$ yr) columns of Tab. 2.
As expected visually from Fig. 1, taking $T\lesssim 0.5$ yr practically guarantees that CHIME J0630 and GX J1627 would be identifiable in the LISA data stream if $P=P_{\rm orb}$. While hopes for other ULPs are low, there is some call for optimism for DA J1832 and A J1935 if the systems contain heavier stars, allowing them to reach significant SNRs when accounting for their sky locations. For instance, if A J1935 contains a neutron star and a reasonably heavy companion, its well-constrained sky location could allow for a boost to the SNR by a factor $\sim 2$, which would give a final value of $\text{SNR}(4\text{yr})\sim 12$ which would be sufficient to claim a confident detection. Combining data from LISA with Taiji and TianQin could also lead to non-negligible boosts in SNR for this and other sources, perhaps enabling detections in marginal scenarios like a WDWD for DA J1832.
Table 2: SNRs for LISA for the systems listed in Tab. 1 for $T=0.5$ yr (middle column) or $T=4$ yr (right) in either the least (largest distances and smallest chirp masses) or most (bracketed values) optimistic scenarios.
Source name SNR (0.5 yr; max) SNR (4 yr; max)
CHIME J0630+25 84.0 (233) 238 (660)
GLEAM-X J1627 31.0 (221.7) 87.8 (627)
GPM J1839–10 0.27 (0.58) 0.77 (1.65)
DA J1832 0.42 (4.33) 1.18 (12.3)
ASKAP J1935 0.22 (2.03) 0.63 (5.73)
GCRT J1745–3009 0.05 (0.50) 0.14 (1.41)
ILT J1101+5521 0.13 (0.22) 0.37 (0.62)
GLEAM-X J0704 0.05 (0.09) 0.15 (0.25)
ASKAP J1839 $<10^{-3}$ (0.004) $<10^{-3}$ (0.01)
4 Conclusions
ULPs remain mysterious, with some studies favouring isolated neutron stars and others favouring WDMDs or compact binaries (see, e.g., Rea et al., 2022; Coti Zelati & Borghese, 2024). We have argued here that if the pulsation period in ULPs matches that of the orbit or close to it, at least a few should be visible to LISA and its sister spacecraft (Tab. 2). If electromagnetic data remain inconclusive, GW observations could help unveil their nature. Importantly, among the known ULPs, those that may be brightest in the LISA band correspond to objects with uncertain classification (Fig. 1).
While some implications of (non-)detections were highlighted in Secs. 2 and 3, a wealth of physical processes may operate in ULPs. For instance, linking radio luminosities to electromotive losses implies a minimum strength for the primary’s magnetic field. Assuming a maser-like emission mechanism, some systems appear to require GigaGauss strengths (Qu & Zhang, 2025). While dynamo action may amplify an internal field (Schreiber et al., 2021), the magnetospheric field may not reflect the interior strength because magnetic burial – a process where field lines are equatorially advected by infalling plasma – will reduce the global dipole moment during accretion while creating strong multipoles (Suvorov & Melatos, 2020). Typical accretion rates in tight-binary polars are $\sim 10^{-10}M_{\odot}\text{ yr}^{-1}$ (Pala et al., 2020) implying that, for a C/O dwarf with $M_{\star}\sim 0.8M_{\odot}$, the accretion timescale matches that of Ohmic diffusion ($\tau_{\Omega}$) at depths corresponding to densities of $\sim 10^{5}\text{ g cm}^{-3}$ where Cumming (2002) estimates $\tau_{\Omega}\lesssim 1$ Gyr. This suggests a timeline for radio activation post detachment if the interaction zone lies at a polar latitude, with the reverse applying near the equator if magnetic flux is compressed there. Resistive relaxation simulations would help elucidate such interactions and their connections with diamagnetic screening, crystallisation, and convection.
Acknowledgements
Support provided by the Conselleria d’Educació, Cultura, Universitats i Ocupació de la Generalitat Valenciana through Prometeo Project CIPROM/2022/13 is gratefully acknowledged. CD acknowledges the Ministerio de Ciencia, Innovación y Universidades, co-funded by the Agencia Estatal de Investigación, the Unión Europea (FSE+), and the Universidad de Alicante, through fellowship JDC2023-052227-I funded by MCIU/AEI/10.13039/501100011033 and the FSE+.